# Core equality of real sequences

### v1

@@ Abstract.
for all bounded sequences $\bm{x}$, provided that $\mathcal{J}$ is a countably generated ideal on $\omega$ and $A$ maps bounded sequences into bounded sequences. Such characterization fails if both $\mathcal{I}$ and $\mathcal{J}$ are the ideal of asymptotic density zero sets. Next, we show that such equality is possible for distinct ideals $\mathcal{I},\mathcal{J}$, answering an open question in [J. Math. Anal. Appl. 321 (2006), 515–523]. Lastly, we prove that, if $\mathcal{J}=\mathrm{Fin}$, the above equality holds for some matrix $A$ if and only if $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}=\mathrm{Fin}\oplus\mathcal{P}(\omega)$.
@@ 1. Introduction
Let $\mathcal{I}$ be an ideal on the natural numbers $\omega$, that is, a family of subsets of $\omega$ closed under subsets and finite unions. Unless otherwise stated, it is assumed that $\mathcal{I}$ contains the family $\mathrm{Fin}$ of finite sets and that $\omega\notin\mathcal{I}$. We denote by $\mathcal{I}^{+}:=\mathcal{P}(\omega)\setminus\mathcal{I}$ and :=$\mathcal{I}^{\star}:=\{S\subseteq\omega:\omega\setminus S\in\mathcal{I}\}$ the family of $\mathcal{I}$-positive sets and the dual filter of $\mathcal{I}$, respectively. Ideals are regarded as subsets of the Cantor space $\{0,1\}^{\omega}$, hence it is possible to speak about $F_{\sigma}$-ideals, analytic ideals, meager ideals, etc. An important example of ideal is the family $\mathcal{Z}$ of sets $S\subseteq\omega$ with asymptotic density zero, that is, $|S\cap[0,n]|=o(n)$ as $n\to\infty$.
@@
where $\mathrm{co}$ denotes the convex hull operator and $\overline{\mathrm{co}}$ its closure, see [14, 21]. In the cases where $\mathcal{I}=\mathrm{Fin}$ and $\mathcal{I}=\mathcal{Z}$, the $\mathcal{I}$-core of $\bm{x}$ is usually called “Knopp core”  and “statistical core,”  respectively, see [12, 13, 18, 28]. Let also $\Gamma_{\bm{x}}(\mathcal{I})$ denote the set of $\mathcal{I}$-cluster points of $\bm{x}$, that is, the set of all $\eta\in X$ such that $\{n\in\omega:x_{n}\in U\}\in\mathcal{I}^{+}$ for all neighborhoods $U$ of $\eta$. It has been shown in [21, Corollary 2.3] that, if $\bm{x}$ is bounded real sequence, then $\mathrm{core}_{\bm{x}}(\mathcal{I})=\mathrm{co}(\Gamma_{\bm{x}}(\mathcal{I}))$. In addition, since $\Gamma_{\bm{x}}(\mathcal{I})$ is a nonempty compact set, see e.g. [24, Lemma 3.1], it follows that
$\mathrm{core}_{\bm{x}}(\mathcal{I})=\left[\,\mathcal{I}\text{-}\liminf\bm{x},\mathcal{I}\text{-}\limsup\bm{x}\,\right]$
@@
Given an infinite matrix $A=(a_{n,k}:n,k\in\omega)$, we denote by $\mathrm{dom}(A)$ its domain, that is, the set of real sequences $\bm{x}=(x_{n})$ which are $A$-summable, namely, the $A$-transformed sequences $A\bm{x}:=\left(A_{n}\bm{x}:n\in\omega\right)$ are well defined, where
$A_{n}\bm{x}:=\sum_{k\in\omega}a_{n,k}x_{k}\quad\text{ for all }n\in\omega,$
provided that each series above is convergent. We write $A\geq 0$ if $a_{n,k}\geq 0$ for all $n,k\in\omega$. Given sequence spaces $\mathcal{A},\mathcal{B}\subseteq\mathbf{R}^{\omega}$, we denote by $(\mathcal{A},\mathcal{B})$ the family of infinite matrices $A$ such that $\mathcal{A}\subseteq\mathrm{dom}(A)$ and $A\bm{x}\in\mathcal{B}$ for all $\bm{x}\in\mathcal{A}$. For instance, it is well known that $A\in(\ell_{\infty},\ell_{\infty})$ if and only if $A\in(c_{0},\ell_{\infty})$ if and only if $\|A\|<\infty$, where $\|A\|:=\sup_{n}\sum_{k}|a_{n,k}|$, see e.g. [5, Theorem 2.3.5]. Here, as usual, $\ell_{\infty}$ and $c$ stand for the vector space of bounded real sequences and convergent real sequences, respectively. $\ell_{\infty}$ and all its subspaces are endowed with the topology induced by the supremum norm.
@@
$A\in(c_{b}(\mathcal{I}),c_{b}(\mathcal{J})\quad\text{ and }\quad\mathcal{I}\text{-}\lim\bm{x}=\mathcal{J}\text{-}\lim A\bm{x}\,\,\text{ for all }\bm{x}\in c_{b}(\mathcal{I}),$
@@
In the statement above, recall that an ideal $\mathcal{J}$ on $\omega$ is countably generated if there exists a sequence $(Q_{k}:k\in\omega)$ of subsets of $\omega$ such that $S\in\mathcal{I}$ if and only if $S\subseteq\bigcup_{k\in F}Q_{k}$ for some finite $F\in\mathrm{Fin}$. Examples of countably generated ideals include $\mathrm{Fin}$ and the isomorphic copies of $\mathrm{Fin}\times\emptyset:=\{S\subseteq\omega^{2}:\exists n\in\omega,S\subseteq[0,n]\times\omega\}$ and $\mathrm{Fin}\oplus\mathcal{P}(\omega):=\{S\subseteq\{0,1\}\times\omega:|S\cap(\{0\}\times\omega)|<\infty\}$, cf. [19, Remark 2.16].
The above result extends the classical Silverman–Toeplitz characterization, which corresponds to the case $\mathcal{I}=\mathcal{J}=\mathrm{Fin}$. Lastly, it is worth to remark that Theorem 1.1 does not hold for arbitrarily ideals: indeed, there exists a $(\mathcal{Z},\mathcal{Z})$-regular which does not satisfy condition (T3), see [7, Theorem 1.4].
@@ 2. Main results
There exists an infinite real matrix $A\in(\ell_{\infty},\ell_{\infty})$ satisfying equality (1) with $\mathcal{J}=\mathrm{Fin}$ if and only if $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}=\mathrm{Fin}\oplus\mathcal{P}(\omega)$.
The latter result is also related to the question posed by Mazur in The Scottish Book whether the notion of statistical convergence (i.e. $\mathcal{Z}$-convergence) of bounded sequences is equivalent to some matrix summability method, see [10] and references therein. A positive answer has been given by Khan and Orhan in [16, Theorem 2.2].
Based on the previous observations, we leave as open question to check whether there exists a matrix $A\in(\ell_{\infty},\ell_{\infty})$ satisfying equality (1) with $\mathcal{I}=\mathrm{Fin}$ if and only if $\mathcal{J}$ is meager. On the same direction, it would be interesting to know if the condition $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ is also necessary in the statament of Theorem 2.4.
@@ 3. Proofs
Let $\bm{x}$ and $\bm{y}$ be two relatively compact sequences taking values in a locally convex topological vector space. Let $\mathcal{I}$ be an ideal on $\omega$. Then
@@
It follows by [24, Lemma 3.5] that $\Gamma_{\bm{x}}(\mathcal{I})=\Gamma_{\bm{y}}(\mathcal{I})$. We conclude by [21, Theorem 2.2] and [14, Theorem 3.4] that $\mathrm{core}_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{y}}(\mathcal{I})=\mathrm{core}_{\bm{y}}(\mathcal{I})$. ∎
@@
Only If part. Pick a sequence $\bm{x}\in c_{b}(\mathcal{I})$ and define $\eta:=\mathcal{I}\text{-}\lim\bm{x}$, so that $\Gamma_{\bm{x}}(\mathcal{I})=\{\eta\}$. It follows that $A\bm{x}$ is well-defined bounded sequence and, thanks to equality (1), that $\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I})=\{\eta\}$. We conclude by [21, Proposition 4.2] that $\mathcal{J}\text{-}\lim A\bm{x}=\eta$, therefore $A$ is $(\mathcal{I},\mathcal{J})$-regular.
It follows by Theorem 1.1 that $\mathcal{J}\text{-}\lim_{n}|a_{n,k}|=0$ for all $k\in\omega$. At this point, pick a set $E\in\mathcal{I}^{+}$. Thanks to [19, Corollary 4.3], there exists a $\{-1,0,1\}$-valued sequence $\bm{x}$ supported on $E$ such that
@@
Define $F:=\{n\in E:x_{n}=1\}$ and $G:=\{n\in E:x_{n}=-1\}$, so that $\{F,G\}$ is a partition of $E$ and $\bm{x}:=\bm{1}_{F}-\bm{1}_{G}$ (hereafter, $\bm{1}_{S}$ stands for the characteristic function of $S$). Note that, since $\mathcal{I}$ is an ideal, then at least one between $F$ and $G$ is an $\mathcal{I}$-positive set. For convenience, let $h:\mathbf{R}\to\mathbf{R}$ be the function defined by $\eta\mapsto|\eta|$. Since $h$ is continuous, it follows by [23, Proposition 3.2] that
@@
where $h(A\bm{x})=(h(A_{n}\bm{x}):n\in\omega)$. Taking into account equality (1) and that the sequence $\bm{x}=\bm{1}_{F}-\bm{1}_{G}$ has at least a $\mathcal{J}$-cluster point in $\{-1,1\}$, we conclude that
$\begin{split}\mathcal{J}\text{-}\limsup_{n\to\infty}\sum_{k\in E}|a_{n,k}|&=\max\,\Gamma_{h(A\bm{x})}(\mathcal{J})=\max\,h\left[\Gamma_{A\bm{x}}(\mathcal{J})\right]=\max\,h\left[\Gamma_{\bm{x}}(\mathcal{J})\right]=1.\end{split}$
@@
$\delta:=\min\left\{\,\frac{\varepsilon}{2+\eta+4\|\bm{x}\|},\,1\right\}\quad\text{ and }\quad E:=\left\{k\in\omega:x_{k}\geq\eta-\delta\right\}.$
@@
$\begin{split}A_{n}\bm{x}&=\sum_{k\in E}a_{n,k}^{+}x_{k}+\sum_{k\in E^{c}}a_{n,k}^{+}x_{k}-\sum_{k\in\omega}a_{n,k}^{-}x_{k}\\ &\geq(1-\delta)(\eta-\delta)-2\delta\|\bm{x}\|-2\delta\|\bm{x}\|\\ &\geq\eta-\delta(1+\eta+4\|\bm{x}\|)\geq\eta-\varepsilon.\end{split}$
@@
$\begin{split}A_{n}\bm{x}&=\sum_{k\in E}a_{n,k}^{+}x_{k}+\sum_{k\in E^{c}}a_{n,k}^{+}x_{k}-\sum_{k\in\omega}a_{n,k}^{-}x_{k}\\ &\leq(1+\delta)(\eta+\delta)+2\delta\|\bm{x}\|+2\delta\|\bm{x}\|\\ &\leq\eta+\delta(2+\eta+4\|\bm{x}\|)\leq\eta+\varepsilon.\end{split}$
@@
This implies that $\Gamma_{\bm{x}}(\mathcal{I})=\Gamma_{A\bm{x}}(\mathcal{J})$, therefore equality (1) holds. ∎
@@
    a Fréchet ideal if, for every $S\in\mathcal{I}^{+}$, the ideal $\mathcal{I}\upharpoonright S:=\mathcal{I}\cap\mathcal{P}(S)$ is not tall.
@@
If part. If $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}=\mathrm{Fin}\oplus\mathcal{P}(\omega)$ then $\mathcal{I}\leq_{\mathrm{RK}}\mathrm{Fin}$. Indeed, in both cases, one can choose the witnessing map $h:\omega\to\omega$ defined by $h(n)=\lfloor n/2\rfloor$ for all $n\in\omega$. The claim follows by Theorem 2.4.
@@
Let us suppose for the sake of contradiction that $\mathcal{I}$ is not a $P^{+}$-ideal, hence it is possible to fix a strictly decreasing sequence $(I_{n}:n\in\omega)$ in $\mathcal{I}$ such that for all sequences $(F_{n}:n\in\omega)$ of finite sets with $F_{n}\subseteq I_{n}\setminus I_{n+1}$ we have $\bigcup_{n}F_{n}\in\mathcal{I}$. Since $I_{0}\in\mathcal{I}^{+}$ it follows by equality (1) that lim $\limsup A\bm{1}_{I_{0}}=1$. Hence there exists $n_{0}\in\omega$ such that $A_{n_{0}}\bm{1}_{I_{0}}>1-2^{-0}$. Set $p_{0}:=0$ and pick an integer $q_{0}>p_{0}$ such that $\sum_{k\in F_{0}}a_{n_{0},k}\bm{1}_{I_{0}}(k)>1-2^{-0}$, where $F_{0}:=\omega\cap[p_{0},q_{0}]$. Recall also that $A$ is $(\mathcal{I},\mathrm{Fin})$-regular since it satisfies (1), hence it is regular. In particular, by Theorem 1.1, we have $\lim_{k}a_{n,k}=0$ and $\sum_{k}|a_{n,k}|<\infty$ for all $n,k\in\omega$. At this point, suppose that $n_{i-1}$ and $F_{i-1}:=\omega\cap[p_{i-1},q_{i-1}]$ have been defined for some positive integer $i$. Then, proceed recursively as follows:
@@
It is easy to see that, if $S$ is finite, then $(I_{n})$ cannot be a sequence which witnesses the non $P$-property of $\mathcal{I}$: indeed, in such case, $I:=I_{0}$ if $S=\emptyset$ or $I:=I_{1+\max S}$ if $S\neq\emptyset$ satisfies $I_{n}\setminus I\in\mathrm{Fin}$ for all $n\in\omega$. Hence $S$ has to be infinite, which implies that, passing to a suitable subsequence, we can assume without loss of generality that $D_{n}$ is infinite for all $n\in\omega$.
@@
Thus =$\mathcal{I}=\{S\subseteq\omega:\mathcal{I}\text{-}\lim\bm{1}_{S}=0\}=\{S\subseteq\omega:\lim A^{+}\bm{1}_{S}=0\}=\{S\subseteq\omega:\mu^{\star}(S)=0\}$. Since $\mathcal{I}$ is not tall, it follows by [8, Proposition 7.2] that $\lim_{n}\sup_{k}a_{n,k}^{+}=0$. In addition, recalling that $A^{+}$ is a nonnegative matrix, we have also that $\lim_{n}a_{n,k}^{+}=0$ and $\sum_{k}a_{n,k}^{+}<\infty$ for all $n,k\in\omega$ by Theorem 1.1. It follows by [8, Theorem 6.2] that the function $\mu^{\star}$ has the weak Darboux property, i.e., for each $S\subseteq\omega$ and $y\in[0,\mu^{\star}(S)]$ there exists $X\subseteq S$ such that $\mu^{\star}(X)=y$, cf. [26, Section 2]. This implies that there exists a decreasing sequence $(I_{m}:m\in\omega)$ of subsets of $\omega$ such that
@@
$\mathcal{I}$ is a Fréchet ideal.

### v2

@@ Abstract.
for all bounded sequences $\bm{x}$, provided that $\mathcal{J}$ is a countably generated ideal on $\omega$ and $A$ maps bounded sequences into bounded sequences. Such characterization fails if both $\mathcal{I}$ and $\mathcal{J}$ are the ideal of asymptotic density zero sets. Next, we show that such equality is possible for distinct ideals $\mathcal{I},\mathcal{J}$, answering an open question in [J. Math. Anal. Appl. 321 (2006), 515–523]. Lastly, we prove that, if $\mathcal{J}=\mathrm{Fin}$, the above equality holds for some matrix $A$ if and only if $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ on $\omega$.
@@ 1. Introduction
Let $\mathcal{I}$ be an ideal on the natural numbers $\omega$, that is, a family of subsets of $\omega$ closed under subsets and finite unions. Unless otherwise stated, it is assumed that $\mathcal{I}$ contains the family $\mathrm{Fin}$ of finite sets and that $\omega\notin\mathcal{I}$. We denote by $\mathcal{I}^{+}:=\mathcal{P}(\omega)\setminus\mathcal{I}$ and :=$\mathcal{I}^{\star}:=\{S\subseteq\omega:\omega\setminus S\in\mathcal{I}\}$ the family of $\mathcal{I}$-positive sets and the dual filter of $\mathcal{I}$, respectively. Ideals are regarded as subsets of the Cantor space $\{0,1\}^{\omega}$, hence it is possible to speak about $F_{\sigma}$-ideals, analytic ideals, meager ideals, etc. An important example of an ideal is the family $\mathcal{Z}$ of sets $S\subseteq\omega$ with asymptotic density zero, that is, $|S\cap[0,n]|=o(n)$ as $n\to\infty$.
@@
where $\mathrm{co}$ denotes the convex hull operator and $\overline{\mathrm{co}}$ its closure, see [14, 22]. In the cases where $\mathcal{I}=\mathrm{Fin}$ and $\mathcal{I}=\mathcal{Z}$, the $\mathcal{I}$-core of $\bm{x}$ is usually called “Knopp core”  and “statistical core,”  respectively, see [12, 13, 18, 29]. Let also $\Gamma_{\bm{x}}(\mathcal{I})$ denote the set of $\mathcal{I}$-cluster points of $\bm{x}$, that is, the set of all $\eta\in X$ such that $\{n\in\omega:x_{n}\in U\}\in\mathcal{I}^{+}$ for all neighborhoods $U$ of $\eta$. It has been shown in [22, Corollary 2.3] that, if $\bm{x}$ is bounded real sequence, then $\mathrm{core}_{\bm{x}}(\mathcal{I})=\mathrm{co}(\Gamma_{\bm{x}}(\mathcal{I}))$. In addition, since $\Gamma_{\bm{x}}(\mathcal{I})$ is a nonempty compact set, see e.g. [25, Lemma 3.1], it follows that
$\mathrm{core}_{\bm{x}}(\mathcal{I})=\left[\,\mathcal{I}\text{-}\liminf\bm{x},\mathcal{I}\text{-}\limsup\bm{x}\,\right]$ (1)
@@
Given an infinite matrix $A=(a_{n,k}:n,k\in\omega)$, we denote by $\mathrm{dom}(A)$ its domain, that is, the set of real sequences $\bm{x}=(x_{n})$ for which the $A$-transformed sequences $A\bm{x}:=\left(A_{n}\bm{x}:n\in\omega\right)$ are well defined, which means that the series
$A_{n}\bm{x}:=\sum_{k\in\omega}a_{n,k}x_{k}$
is convergent to a (finite) real number for each $n\in\omega$. We write $A\geq 0$ if $a_{n,k}\geq 0$ for all $n,k\in\omega$. Given sequence spaces $\mathcal{A},\mathcal{B}\subseteq\mathbf{R}^{\omega}$, we denote by $(\mathcal{A},\mathcal{B})$ the family of infinite matrices $A$ such that $\mathcal{A}\subseteq\mathrm{dom}(A)$ and $A\bm{x}\in\mathcal{B}$ for all $\bm{x}\in\mathcal{A}$. For instance, it is well known that $A\in(\ell_{\infty},\ell_{\infty})$ if and only if $A\in(c,c)$ if and only if $\|A\|<\infty$, where $\|A\|:=\sup_{n}\sum_{k}|a_{n,k}|$, see e.g. [5, Theorem 2.3.5]. Here, as usual, $\ell_{\infty}$ and $c$ stand for the vector space of bounded real sequences and convergent real sequences, respectively. $\ell_{\infty}$ and all its subspaces are endowed with the topology induced by the supremum norm.
@@
$A\in(c_{b}(\mathcal{I}),c_{b}(\mathcal{J}))\quad\text{ and }\quad\mathcal{I}\text{-}\lim\bm{x}=\mathcal{J}\text{-}\lim A\bm{x}\,\,\text{ for all }\bm{x}\in c_{b}(\mathcal{I}),$
@@
In the statement above, recall that an ideal $\mathcal{J}$ on $\omega$ is countably generated if there exists a sequence $(Q_{k}:k\in\omega)$ of subsets of $\omega$ such that $S\in\mathcal{I}$ if and only if $S\subseteq\bigcup_{k\in F}Q_{k}$ for some finite $F\in\mathrm{Fin}$. Examples of countably generated ideals include $\mathrm{Fin}$ and the isomorphic copies on $\omega$ of $\mathrm{Fin}\times\{\emptyset\}:=\{S\subseteq\omega^{2}:\exists n\in\omega,S\subseteq[0,n]\times\omega\}$ and $\mathrm{Fin}\oplus\mathcal{P}(\omega):=\{S\subseteq\{0,1\}\times\omega:|S\cap(\{0\}\times\omega)|<\infty\}$, cf. [19, Remark 2.16]. Hereafter, an ideal $\mathcal{I}$ on a countably infinite set $W$ is said to be an isomorphic copy of an ideal $\mathcal{J}$ on $\omega$ if there exists a bijection $f:W\to\omega$ such that $S\in\mathcal{J}$ if and only if $f^{-1}[S]\in\mathcal{I}$ for each $S\subseteq\omega$.)
The above result extends the classical Silverman–Toeplitz characterization, which corresponds to the case $\mathcal{I}=\mathcal{J}=\mathrm{Fin}$. Lastly, it is worth mentioning that Theorem 1.1 does not hold for arbitrarily ideals: indeed, there exists a $(\mathcal{Z},\mathcal{Z})$-regular which does not satisfy condition (T3), see [7, Theorem 1.4].
@@ 2. Main results
There exists an infinite real matrix $A\in(\ell_{\infty},\ell_{\infty})$ satisfying equality (2) with $\mathcal{J}=\mathrm{Fin}$ if and only if $\mathcal{I}=\mathrm{Fin}$ or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ on $\omega$.
(Equivalently, the latter condition means =$\mathcal{I}=\{S\subseteq\omega:S\cap T\in\mathrm{Fin}\}$ for some infinite $T\subseteq\omega$.) It is worth noting that Theorem 2.5 is also related to the question posed by Mazur in The Scottish Book whether the notion of statistical convergence (i.e., $\mathcal{Z}$-convergence) of bounded sequences is equivalent to some matrix summability method, see [10] and references therein. A positive answer has been given by Khan and Orhan in [16, Theorem 2.2].
Based on the previous observations, we leave as an open question to check whether there exists a matrix $A\in(\ell_{\infty},\ell_{\infty})$ satisfying equality (2) with $\mathcal{I}=\mathrm{Fin}$ if and only if $\mathcal{J}$ is meager. In the same direction, it would be interesting to know if the condition $\mathcal{I}\leq_{\mathrm{RK}}\mathcal{J}$ is also necessary in the statament of Theorem 2.4.
@@ 3. Proofs
Let $\bm{x}$ and $\bm{y}$ be two relatively compact sequences taking values in a locally convex topological vector space $X$. Let $\mathcal{I}$ be an ideal on $\omega$. Then
@@
Since $X$ is, in particular, a topological group, it follows by [25, Lemma 3.5] and the hypothesis -$\mathcal{I}\text{-}\lim(\bm{x}-\bm{y})=0$ that $\Gamma_{\bm{x}}(\mathcal{I})=\Gamma_{\bm{y}}(\mathcal{I})$. Now, let $K_{\bm{x}}$ be the closure of the image $\{x_{n}:n\in\omega\}$, which is a compact subset of the locally convex space $X$. Of course, this implies that $\{x_{n}:n\in\omega\}\subseteq K_{\bm{x}}+U$ for every open neighborhood $U$ of 00 (which means that $\bm{x}$ is “$\mathcal{I}^{\star}$-asymptotically $K_{\bm{x}}$-controlled,” using the terminology in [14, Definition 3.2]). It follows by [14, Theorem 3.4] that $\mathrm{core}_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{x}}(\mathcal{I})$; cf. also [22, Theorem 2.2] in the case of first countable locally convex spaces. With an analog reasoning on the sequence $\bm{y}$, we conclude that $\mathrm{core}_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{x}}(\mathcal{I})=\overline{\mathrm{co}}\,\Gamma_{\bm{y}}(\mathcal{I})=\mathrm{core}_{\bm{y}}(\mathcal{I})$. ∎
@@
Only If part. Pick a sequence $\bm{x}\in c_{b}(\mathcal{I})$ and define $\eta:=\mathcal{I}\text{-}\lim\bm{x}$, so that $\Gamma_{\bm{x}}(\mathcal{I})=\{\eta\}$. It follows that $A\bm{x}$ is well-defined bounded sequence and, thanks to equality (2), that $\mathrm{core}_{A\bm{x}}(\mathcal{J})=\mathrm{core}_{\bm{x}}(\mathcal{I})=\{\eta\}$. We conclude by [22, Proposition 4.2] that $\mathcal{J}\text{-}\lim A\bm{x}=\eta$, therefore $A$ is $(\mathcal{I},\mathcal{J})$-regular.
At this point, pick a set $E\in\mathcal{I}^{+}$. Recall that the hypothesis $A\in(\ell_{\infty},\ell_{\infty})$ is equivalent to $\sup_{n\in\omega}\sum_{k}|a_{n,k}|<\infty$, see e.g. [5, Theorem 2.3.5]. In addition, it follows by Theorem 1.1 that $\mathcal{J}\text{-}\lim_{n}|a_{n,k}|=0$ for all $k\in\omega$. This implies that conditions (K1)–(K3) in the statement of [19, Corollary 4.3] are satisfied (in the case where $d=m=1$, so that $a_{n,k}(i,j)$ is simply $a_{n,k}$ considering that both $i$ and $j$ can take only one value). Hence, thanks to [19, Corollary 4.3], there exists a $\{-1,0,1\}$-valued sequence $\bm{x}$ supported on $E$ such that
@@
(Again, equation (3) is a rewriting of the claim in [19, Corollary 4.3] taking into account that both variables $i$ and $j$ can take only value.) Define $F:=\{n\in E:x_{n}=1\}$ and $G:=\{n\in E:x_{n}=-1\}$, so that $\{F,G\}$ is a partition of $E$ and $\bm{x}:=\bm{1}_{F}-\bm{1}_{G}$ (hereafter, $\bm{1}_{S}$ stands for the characteristic function of $S$). Note that, since $\mathcal{I}$ is an ideal, then at least one between $F$ and $G$ is an $\mathcal{I}$-positive set. For convenience, let $h:\mathbf{R}\to\mathbf{R}$ be the function defined by $\eta\mapsto|\eta|$. Since $h$ is continuous, it follows by [24, Proposition 3.2] that
@@
[…]
$\begin{split}\mathcal{J}\text{-}\limsup_{n\to\infty}\sum_{k\in E}|a_{n,k}|&=\max\,\Gamma_{h(A\bm{x})}(\mathcal{J})=\max\,h\left[\Gamma_{A\bm{x}}(\mathcal{J})\right]\\ &=\max\{|\mathcal{J}\text{-}\limsup A\bm{x}|,|\mathcal{J}\text{-}\liminf A\bm{x}|\}\\ &=\max\{|\mathcal{J}\text{-}\limsup\bm{x}|,|\mathcal{J}\text{-}\liminf\bm{x}|\}=\max\,h\left[\Gamma_{\bm{x}}(\mathcal{J})\right]=1.\end{split}$
@@
$\delta:=\min\left\{\,\frac{\varepsilon}{2+\eta+4\|x\|},\,1\right\}\quad\text{ and }\quad E:=\left\{k\in\omega:x_{k}\geq\eta-\delta\right\}.$
@@
$\begin{split}A_{n}\bm{x}&=\sum_{k\in E}a_{n,k}^{+}x_{k}+\sum_{k\in E^{c}}a_{n,k}^{+}x_{k}-\sum_{k\in\omega}a_{n,k}^{-}x_{k}\\ &\geq(1-\delta)(\eta-\delta)-2\delta\|x\|-2\delta\|x\|\\ &\geq\eta-\delta(1+\eta+4\|x\|)\geq\eta-\varepsilon.\end{split}$
@@
$\begin{split}A_{n}\bm{x}&=\sum_{k\in E}a_{n,k}^{+}x_{k}+\sum_{k\in E^{c}}a_{n,k}^{+}x_{k}-\sum_{k\in\omega}a_{n,k}^{-}x_{k}\\ &\leq(1+\delta)(\eta+\delta)+2\delta\|x\|+2\delta\|x\|\\ &\leq\eta+\delta(2+\eta+4\|x\|)\leq\eta+\varepsilon.\end{split}$
@@
This implies that $\Gamma_{\bm{x}}(\mathcal{I})=\Gamma_{A\bm{x}}(\mathcal{J})$, so that by [22, Corollary 2.3] we get
$\mathrm{core}_{\bm{x}}(\mathcal{I})=\mathrm{co}(\Gamma_{\bm{x}}(\mathcal{I}))=\mathrm{co}(\Gamma_{A\bm{x}}(\mathcal{J}))=\mathrm{core}_{A\bm{x}}(\mathcal{J}).$
Therefore equality (2) holds. ∎
@@
    nowhere tall if, for every $S\in\mathcal{I}^{+}$, the ideal $\mathcal{I}\upharpoonright S:=\mathcal{I}\cap\mathcal{P}(S)$ is not tall.
@@
If part. Suppose that =$\mathcal{I}=\{S\subseteq\omega:S\cap T\in\mathrm{Fin}\}$ for some infinite $T\subseteq\omega$ (i.e., $\mathcal{I}=\mathrm{Fin}$ if $T$ is cofinite or $\mathcal{I}$ is an isomorphic copy of $\mathrm{Fin}\oplus\mathcal{P}(\omega)$ on $\omega$ otherwise). Then $\mathcal{I}\leq_{\mathrm{RK}}\mathrm{Fin}$. In fact, if $(t_{n}:n\in\omega)$ denotes the increasing enumeration of elements of $T$, one can choose the witnessing map $h:\omega\to\omega$ defined by $h(n)=t_{n}$ for all $n\in\omega$. The claim follows by Theorem 2.4.
@@
Let us suppose for the sake of contradiction that $\mathcal{I}$ is not a $P^{+}$-ideal, hence it is possible to fix a strictly decreasing sequence $(I_{n}:n\in\omega)$ in $\mathcal{I}^{+}$ such that for all sequences $(F_{n}:n\in\omega)$ of finite sets with $F_{n}\subseteq I_{n}\setminus I_{n+1}$ we have $\bigcup_{n}F_{n}\in\mathcal{I}$. Since $I_{0}\in\mathcal{I}^{+}$ it follows by equality (2) that lim $\limsup A\bm{1}_{I_{0}}=1$. Hence there exists $n_{0}\in\omega$ such that $A_{n_{0}}\bm{1}_{I_{0}}>1-2^{-0}$. Set $p_{0}:=0$ and pick an integer $q_{0}>p_{0}$ such that $\sum_{k\in F_{0}}a_{n_{0},k}\bm{1}_{I_{0}}(k)>1-2^{-0}$, where $F_{0}:=\omega\cap[p_{0},q_{0}]$. Recall also that $A$ is $(\mathcal{I},\mathrm{Fin})$-regular since it satisfies (2), hence it is regular. In particular, by Theorem 1.1, we have $\lim_{k}a_{n,k}=0$ and $\sum_{k}|a_{n,k}|<\infty$ for all $n,k\in\omega$. At this point, suppose that $n_{i-1}$ and $F_{i-1}:=\omega\cap[p_{i-1},q_{i-1}]$ have been defined for some positive integer $i$. Then, proceed recursively as follows:
@@
It is easy to see that, if $S$ is finite, then $(I_{n}:n\in\omega)$ cannot be a sequence witnessing that $\mathcal{I}$ is not a $P$-ideal: indeed, in such case, $I:=I_{0}$ if $S=\emptyset$ or $I:=I_{1+\max S}$ if $S\neq\emptyset$ satisfies $I_{n}\setminus I\in\mathrm{Fin}$ for all $n\in\omega$. Hence $S$ has to be infinite, which implies that, passing to a suitable subsequence, we can assume without loss of generality that $D_{n}$ is infinite for all $n\in\omega$.
@@
Thus =$\mathcal{I}=\{S\subseteq\omega:\mathcal{I}\text{-}\lim\bm{1}_{S}=0\}=\{S\subseteq\omega:\lim A^{+}\bm{1}_{S}=0\}=\{S\subseteq\omega:\mu^{\star}(S)=0\}$. Since $\mathcal{I}$ is not tall, it follows by [8, Proposition 7.2] that $\lim_{n}\sup_{k}a_{n,k}^{+}=0$. In addition, recalling that $A^{+}$ is a nonnegative matrix, we have also that $\lim_{n}a_{n,k}^{+}=0$ and $\sum_{k}a_{n,k}^{+}<\infty$ for all $n,k\in\omega$ by Theorem 1.1. It follows by [8, Theorem 6.2] that the function $\mu^{\star}$ has the weak Darboux property, i.e., for each $S\subseteq\omega$ and $y\in[0,\mu^{\star}(S)]$ there exists $X\subseteq S$ such that $\mu^{\star}(X)=y$, cf. [27, Section 2]. This implies that there exists a decreasing sequence $(I_{m}:m\in\omega)$ of subsets of $\omega$ such that
@@
$\mathcal{I}$ is a nowhere tall ideal.